# K–P Quantum Neural Networks

### v1

@@ Abstract
We present an extension of K–P time-optimal quantum control solutions using global Cartan $KAK$ decompositions for geodesic-based solutions. Extending recent time-optimal constant–$\theta$ control results, we integrate Cartan methods into equivariant quantum neural network (EQNN) for quantum control tasks. We show that a finite-depth limited EQNN ansatz equipped with Cartan layers can replicate the constant–$\theta$ sub-Riemannian geodesics for K–P problems. We demonstrate how for certain classes of control problem on Riemannian symmetric spaces, gradient-based training using an appropriate cost function converges to certain global time-optimal solutions when satisfying simple regularity conditions. This generalises prior geometric control theory methods and clarifies how optimal geodesic estimation can be performed in quantum machine learning contexts.
@@ 1 Introduction
Time-optimal control of quantum systems is central to many areas of quantum technology, ranging from fast gate synthesis in quantum computing to high-fidelity pulse shaping in nuclear magnetic resonance [7, 12, 5]. The K–P problem [4] is a canonical formulation of such time-optimal tasks. K–P problems involve a semisimple Lie algebra $\mathfrak{g}$ into $\mathfrak{k}\oplus\mathfrak{p}$ under a Cartan (or involution-based) decomposition [8]. The physically available (horizontal) controls come from $\mathfrak{p}$, while the compact part $\mathfrak{k}=[\mathfrak{p},\mathfrak{p}]$ must be generated indirectly via commutators. Previous work has shown that sub-Riemannian geometry on $\mathfrak{g}$ yields geodesics for locally time-optimal motion [6, 2]. Recently, a new method of optimal control was demonstrated [14] employing global Cartan $KAK$ decompositions and enforcing a constant–$\theta$ condition, one obtains an analytically solvable geodesic for quantum control problems on certain classes of Riemannian symmetric space. Other recent work has examined relaxing the requirement of full equivariance for qubit systems using variational quantum algorithms generated by horizontal elements $\mathfrak{p}$ of $KAK$ structures [16]. This partial respecting of symmetry translations is equivalent to classes of sub-Riemannian control problems invariant under translations $\mathfrak{p}$ but not generators of rotations $\mathfrak{k}$. We extend these results in two directions:
@@
    K–P QNN (EQNN) Integration.  We show how to parameterize the same Cartan-based geodesics in a neural-network ansatz that is equivariant to the underlying symmetry group. Extending existing work, we show how as with full EQNNs, composing layer-wise networks using our global method which respects the $\mathfrak{k}/\mathfrak{p}$ partition enables the networks (and importantly, their outputs) to respect sub-Riemannian symmetries. We subsequently explore how K–P respecting networks converge to approximate sub-Riemannian geodesics, without requiring the user to solve the geodesic equations symbolically.
@@ 2.1 Cartan decompositions
$K=\exp(\mathfrak{k})$ is a subgroup of $G$. If $G$ is compact and semisimple then $K$ is typically a maximal torus. A typical quantum control scenario involves Hamiltonians comprising generators in $\mathfrak{p}$, while the evolution generated by the $\mathfrak{k}$ part arises via the commutators in (1). We want to implement a target $U_{\mathrm{target}}\in G$ in minimal time subject to an energy cutoff $\|H(t)\|\leq\Omega$, with $H(t)\in\mathfrak{p}$. In sub-Riemannian geometric terms, $\mathfrak{p}$ defines the horizontal distribution. $\mathfrak{k}$ is the vertical direction that can be reached by curvature forms $[\mathfrak{p},\mathfrak{p}]\subset\mathfrak{k}$ [10]. The corresponding group $KAK$ decomposition is given by $G=K\exp(\mathfrak{a})K$, where $\mathfrak{a}\subset\mathfrak{p}$ is an abelian subalgebra (maximally noncompact Cartan). Elements in $\mathfrak{a}$ typically look like $i\Theta$, with $\Theta$ real diagonal in a suitable representation. Then any $U\in G$ can be written:
@@
[14] showed that the time-optimal solutions $\gamma(t)$ can be simplified if we impose $d\Theta(t)=0$ along the path, the so-called constant–$\theta$ condition. Under mild conditions, this yields a closed-form geodesic and the minimal time $T$ is related to $|\sin(\mathrm{ad}_{\Theta})(\Phi)|$, with $\Phi\in\mathfrak{k}$ in the commutant of $\Theta$. The result in [14] can be expressed as follows.
@@
Let $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ be a Cartan decomposition of a compact semisimple Lie algebra and let $\Theta\in\mathfrak{a}\cap\mathfrak{p}$ be in the noncompact Cartan subalgebra. Suppose $\Phi\in\mathfrak{k}$ commutes with $\Theta$. Then if $H(t)\in\mathfrak{p}$ saturates $\|H\|=\Omega$ and satisfies the minimal connection plus $\dot{\Theta}(t)=0$, the minimum time $T$ optimal path from $U(0)=\exp(\mathrm{i}\Theta)$ to $U(T)=\exp(-\mathrm{i}X)$ ($X\in\mathfrak{k}$ a target) has length
@@ 3 K–P QNNs
EQNNs architect neural networks to respect group symmetries in ways that facilitate task optimisation [11, 15] and have shown success in quantum optimisation tasks. EQNNs have layers designed so that transformations by a group $G$ act consistently on inputs and outputs [9]. The K–P problem has an explicit decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ associated with the involution $\chi$. A natural route to building an QNN respecting K–P structure is to make the layer transformations equivariant with respect to symmetry subgroup $K$ (the subgroup generated by $\mathfrak{k}$). In this formulation, unitary conjugation $k\in K$ transforms the QNN parameters consistent with Eqn. (1) above. In the simplest sense, an K–P layer can be written:
@@
and hence $U(t)$ could be generated purely by $\mathfrak{p}$. We encode such constraints into the QNN layer as follows.
@@
    Initialize $\Theta$ (frozen). First, we choose an element $\Theta\in\mathfrak{a}\cap\mathfrak{p}$ (so it is in the noncompact Cartan subalgebra). In an N-qubit representation, this might be a block diagonal or simple diagonal with real entries.
@@
    Parameterize commutant $\Phi\in\mathrm{Comm}(\Theta)\cap\mathfrak{k}$. Because $\Phi$ commutes with $\Theta$, $\mathrm{Ad}_{\exp(\mathrm{i}\Theta)}(\Phi)=\Phi$.
@@
Because the layer is built from $\{\Theta,\Phi\}$ with $\Theta$ fixed, the QNN is automatically equivariant under transformations in $K$ that fix $\Theta$ (or map it to an isomorphic subalgebra). In practice, the numerical training need not solve the entire geodesic system explicitly, but the final result (provided local minima are avoided) matches the sub-Riemannian solution.
@@ 4 Existence and uniqueness of K–P circuits
The control parameters $\alpha$ may be learnt and optimised according to a typical VQA optimisatoin algorithm. Next, we show that any local optimum of a standard cost function in the EQNN approach must coincide with the global sub-Riemannian geodesic solution.
@@
Theorems 4.1 and 4.2 show that (for certain choices of subgroup $K$ (and thus symmetric space) that K–P QNNs can be made both complete and globally convergent: (i) the finite-depth circuit is guaranteed to express a minimal-time path; and (ii) gradient-based optimisation (with an appropriate cost) can avoid spurious local minima. We briefly illustrate how K–P QNNs with a can discover the same time-optimal solution as an analytic or geometric approach.
@@ 5 Evaluation: $\Lambda$-systems and $SU(3)$
Arrays store these respective matrices. We construct a finite-depth product of exponentials:
@@
In the code, we approximate the sub-Riemannian path length by the sum of the $\ell_{2}$-norms of the horizontal parameters $\alpha^{(\mathfrak{p})}$ encoded in path\_length. This is a simplified version of $\int_{0}^{T}\|H_{\mathfrak{p}}(t)\|\mathrm{d}t$, sufficient to demonstrate the principle of penalizing the magnitude of $\mathfrak{p}$-controls. We apply a straightforward gradient descent (using JAX auto-differentiation). By Theorem 4.2, no spurious local minima exist if $\boldsymbol{\alpha}$ saturates the bracket generation assumptions and $\|H(t)\|\leq\Omega$. Convergence to fidelity $\approx 1$ and small path length indicates that the learned solution reproduces the sub-Riemannian geodesic described in the main text.
Figure 1: K–P QNN Loss function (Eqn. (6) showing convergence between $U(\alpha)$ and $U_{target}$. Convergence shows the K–P QNN learning the optimal unitary parametrised by controls $\alpha$ (reproducing targets in [1] and [14]).
@@ 6 Conclusions and Outlook
We have shown how K–P QNNs, a form of EQNNs can be naturally constructed using the constant–$\theta$ Cartan decomposition approach to the K–P problem. The synergy arises from the geometric consistency: EQNN layers that keep $\Theta$ fixed and use pulses in $\mathfrak{p}$ effectively trace out the same sub-Riemannian geodesic described by [14]. Our results show that local optimality in a typical QNN variational cost function indeed implies global time-optimality for certain classes of quantum control problems on symmetric spaces $G/K$. Limitations of our method include those set out in [14], particularly that time-optimal sequences are found for only certain targets in $\mathfrak{g}$. Future research directions building on this work may consider:
@@
  • extending to non-compact or indefinite metrics (e.g. involving indefinite Killing forms or non-compact groups); or
@@ 6.0.1 Acknowledgements
The author thanks Chris Jackson for his considerable mentorship and discussions on the topic. This work was independently funded by the author.
[…]
The author declares that they have no conflicts of interest.

### v2

@@ Abstract
We present an extension of K–P time-optimal quantum control solutions using global Cartan $KAK$ decompositions for geodesic-based solutions. Extending recent time-optimal constant–$\theta$ control results, we integrate Cartan methods into equivariant quantum neural network (EQNN) for quantum control tasks. We show that a finite-depth limited EQNN ansatz equipped with Cartan layers can replicate the constant–$\theta$ sub-Riemannian geodesics for K–P problems. We demonstrate how, for certain classes of control problem on Riemannian symmetric spaces, gradient-based training using an appropriate cost function converges to global time-optimal solutions when satisfying simple regularity conditions. This generalises prior geometric control theory methods and clarifies how optimal geodesic estimation can be performed in quantum machine learning contexts.
@@ 1 Introduction
Time-optimal control of quantum systems is central to many areas of quantum technology, ranging from fast gate synthesis in quantum computing to high-fidelity pulse shaping in nuclear magnetic resonance [7, 12, 5]. The K–P problem [4] is a canonical formulation of such time-optimal tasks. K–P problems involve a semisimple Lie algebra $\mathfrak{g}$ into $\mathfrak{k}\oplus\mathfrak{p}$ under a Cartan (or involution-based) decomposition [8]. The horizontal controls come from $\mathfrak{p}$, while the compact part $\mathfrak{k}=[\mathfrak{p},\mathfrak{p}]$ must be generated indirectly via commutators. Previous work has shown that sub-Riemannian geometry on $\mathfrak{g}$ yields geodesics for locally time-optimal motion [6, 2]. Recently, a new method of optimal control was demonstrated [14] employing global Cartan $KAK$ decompositions and enforcing a constant–$\theta$ condition, one obtains an analytically solvable geodesic for quantum control problems on certain classes of Riemannian symmetric space. Other recent work has examined relaxing the requirement of full equivariance for qubit systems using variational quantum algorithms generated by horizontal elements $\mathfrak{p}$ of $KAK$ structures [16]. This partial respecting of symmetry translations is equivalent to classes of sub-Riemannian control problems invariant under translations $\mathfrak{p}$ but not generators of rotations $\mathfrak{k}$. We extend these results in two directions:
@@
    K–P QNN (EQNN) Integration.  We show how to parameterise the same Cartan-based geodesics using a neural-network ansatz that is equivariant to the underlying symmetry group. Extending existing work, we show how as with full EQNNs, composing layer-wise networks using our global method which respects the $\mathfrak{k}/\mathfrak{p}$ partition enables the networks (and importantly, their outputs) to respect sub-Riemannian symmetries. We subsequently explore how K–P respecting networks converge to approximate sub-Riemannian geodesics, without requiring the user to solve the geodesic equations symbolically.
@@ 2 Background
2.1 The $K-P$ Problem
The K–P problem in control theory is a specific class of problems that seeks, for spaces $G$ decomposable into symmetric $K$ and anti-symmetric $P$ subspaces, the optimal horizontal control $H(t)\!\in\!\mathfrak{p}$ with $\|H(t)\|\!\leq\!\Omega$ solution that evolves $U(0)=\mathbb{I}$ to a given $U_{\mathrm{target}}\in G$ in minimum time $T$. This problem can equivalently be framed as seeking the minimal sub-Riemannian length $\int_{0}^{T}\|H_{\mathfrak{p}}(t)\|\,\mathrm{d}t$ over all horizontal curves on the symmetric space $G/K$ satisfying identical endpoint constraints. Geodesic solutions on the compact symmetric space $G/K$ of the corresponding Schrödinger equation can be shown to remain within $K\exp(i\Theta)K$ orbits (which are effectively fixed once the Carten element $\Theta$ is chosen).
@@ 2.2 Cartan decompositions
$K=\exp(\mathfrak{k})$ is a subgroup of $G$. If $G$ is compact and semisimple then $K$ is typically a maximal torus. A typical quantum control scenario involves Hamiltonians comprising generators in $\mathfrak{p}$, while the evolution generated by the $\mathfrak{k}$ part arises via the commutators in (1). We want to implement a target $U_{\mathrm{target}}\in G$ in minimal time subject to an energy cutoff $\|H(t)\|\leq\Omega$, with $H(t)\in\mathfrak{p}$. In sub-Riemannian geometric terms, $\mathfrak{p}$ defines the horizontal distribution. $\mathfrak{k}$ is the vertical direction that can be reached by curvature forms $[\mathfrak{p},\mathfrak{p}]\subset\mathfrak{k}$ [10]. The corresponding group $KAK$ decomposition is given by $G=K\exp(\mathfrak{a})K$, where $\mathfrak{a}\subset\mathfrak{p}$ is an abelian subalgebra (maximally non-compact Cartan). Elements in $\mathfrak{a}$ typically look like $i\Theta$, with $\Theta$ real diagonal in a suitable representation. Then any $U\in G$ can be written:
@@
[14] showed that the time-optimal solutions $\gamma(t)$ can be simplified if we impose $d\Theta(t)=0$ along the path, the so-called constant–$\theta$ condition. Under such conditions, this yields a closed-form geodesic and the minimal time $T$ is related to $|\sin(\mathrm{ad}_{\Theta})(\Phi)|$, with $\Phi\in\mathfrak{k}$ in the commutant of $\Theta$. The result in [14] can be expressed as follows. Fixing $\dot{\Theta}=0$ selects the representative in each $K$-orbit that minimises the Cartan energy functional. This can be expressed in terms of the following theorem.
@@
Let $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ be a Cartan decomposition of a compact semisimple Lie algebra and let $\Theta\in\mathfrak{a}\cap\mathfrak{p}$ be in the non-compact Cartan subalgebra. Suppose $\Phi\in\mathfrak{k}$ commutes with $\Theta$. Then if $H(t)\in\mathfrak{p}$ saturates $\|H\|=\Omega$ and satisfies the minimal connection plus $\dot{\Theta}(t)=0$, the minimum time $T$ optimal path from $U(0)=\exp(\mathrm{i}\Theta)$ to $U(T)=\exp(-\mathrm{i}X)$ ($X\in\mathfrak{k}$ a target) has length
@@ 3 K–P QNNs
EQNNs architect neural networks to respect group symmetries in ways that facilitate task optimisation [11, 15] and have shown success in quantum optimisation tasks. EQNNs have layers designed so that transformations by a group $G$ act consistently on inputs and outputs [9]. The K–P problem has an explicit decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ associated with the involution $\chi$. A route to building an QNN respecting K–P structure is to make the layer transformations equivariant with respect to symmetry subgroup $K$ (the subgroup generated by $\mathfrak{k}$). In this formulation, unitary conjugation $k\in K$ transforms the QNN parameters consistent with Eqn. (1) above. In the simplest sense, an K–P layer can be written:
@@
and hence $U(t)$ could be generated purely by $\mathfrak{p}$. We encode such constraints into the QNN layer as follows:
@@
    Initialise $\Theta$. First, we choose an element $\Theta\in\mathfrak{a}\cap\mathfrak{p}$ (so it is in the noncompact Cartan subalgebra). In an $n$-qubit representation, this might be a block diagonal or simple diagonal with real entries.
@@
    Parameterise commutant $\Phi\in\mathrm{Comm}(\Theta)\cap\mathfrak{k}$. Because $\Phi$ commutes with $\Theta$, $\mathrm{Ad}_{\exp(\mathrm{i}\Theta)}(\Phi)=\Phi$.
@@
Because the layer is built from $\{\Theta,\Phi\}$ with $\Theta$ fixed, the QNN is automatically equivariant under transformations in $K$ that fix $\Theta$ (or map it to an isomorphic subalgebra). In practice, the numerical training need not solve the entire geodesic system explicitly, but the final result (provided local minima are avoided) ought to match the sub-Riemannian solution (that is, a global length-minimising horizontal curve).
@@ 4 Existence and uniqueness of K–P circuits
Because the generators generate a dense subgroup of $G$ (by the Cartan decomposition), the Lie-Trotter approximation in principle allows approximation to arbitrary accuracy up to Trotter error $\varepsilon$ using a finite product of $\mathfrak{p}$ and $\mathfrak{k}$ exponentials. The control parameters $\alpha$ may be learnt and optimised according to a typical VQA optimisation algorithm. Next, we show that any local optimum of a standard cost function in the EQNN approach must coincide with the global sub-Riemannian geodesic solution.
@@
Because the optimal Hamiltonian here is normal (otherwise failing to satisfy the maximum principle), abnormal trajectories are excluded. Stationary points with $\mathcal{L}\geq\Omega T$ are precluded by assumptions that $\nabla_{\boldsymbol{\alpha}}C=0$ and $\lambda(\mathcal{L}-\Omega T)=0$ and $\lambda\geq 0$. Theorems 4.1 and 4.2 show that (for certain choices of subgroup $K$ (and thus symmetric space) that K–P QNNs can be made both complete and globally convergent: (i) the finite-depth circuit is guaranteed to express a minimal-time path; and (ii) gradient-based optimisation (with an appropriate cost) can avoid spurious local minima. We briefly illustrate how K–P QNNs with a can discover the same time-optimal solution as an analytic or geometric approach.
@@ 5 Evaluation: $\Lambda$-systems and $SU(3)$
We construct a finite-depth product of exponentials:
@@
In the code, we approximate the sub-Riemannian path length by the sum of the $\ell_{2}$-norms of the horizontal parameters $\alpha^{(\mathfrak{p})}$ encoded in path\_length. This is a simplified version of $\int_{0}^{T}\|H_{\mathfrak{p}}(t)\|\mathrm{d}t$, sufficient to demonstrate the principle of penalizing the magnitude of $\mathfrak{p}$-controls. We apply a straightforward gradient descent (using JAX auto-differentiation). By Theorem 4.2, no spurious local minima exist if $\boldsymbol{\alpha}$ saturates the bracket generation assumptions and $\|H(t)\|\leq\Omega$. Convergence to fidelity $\approx 1$ and small path length indicates that the learned solution reproduces the sub-Riemannian geodesic.
Figure 1: K–P QNN Loss function (Eqn. (6) showing convergence between $U(\alpha)$ and $U_{target}$. Convergence of both terms shows that the K–P QNN learns the optimal unitary parametrised by controls $\alpha$ (reproducing targets in [1] and [14]).
@@ 6 Conclusions and Outlook
We have shown how K–P QNNs, a form of EQNNs can be constructed using the constant–$\theta$ Cartan decomposition approach to the K–P problem. The synergy arises from the geometric consistency: EQNN layers that keep $\Theta$ fixed and use pulses in $\mathfrak{p}$ effectively trace out the same sub-Riemannian geodesic described by [14]. Our results show that local optimality in a typical QNN variational cost function indeed implies global time-optimality for certain classes of quantum control problems on symmetric spaces $G/K$. Limitations of our method include those set out in [14], particularly that time-optimal sequences are found for only certain targets in $\mathfrak{g}$. We have also assumed no complications arising from cut or critical loci. Future research directions building on this work may consider:
@@
  • extending to non-compact or indefinite metrics (e.g. involving indefinite Killing forms or non-compact groups);
  • examining how critical and cut loci (local component Maxwell sets, conjugate points) affect numerically-obtained local minima;
@@ 6.0.1 Acknowledgements
The author thanks Dr Chris Jackson for his considerable mentorship and discussions on the topic. This work was independently funded by the author. The author declares that they have no conflicts of interest.